A 200g of gas was heated from 50°C to around 120°C. The gas expanded in a piston-cylinder apparatus from 2.0L of initial volume to find a volume of 10.0L against an external pressure of 3.0 atm. Calculate for the change in the internal energy. The specific heat of water is 2.19 J/g×°C)


1
Expert's answer
2022-03-13T18:50:22-0400

Solution;

"\Delta u=q+W"

"q=mC\Delta T"

"q=200\u00d72.19(120-50)"

"q=30660J"

"W=-P(V_2-V_1)"

"W=-303.975(10-2)"

"W=2431.8J"

Hence;

"\Delta u=30660-2431.8=28.23kJ"